### v1

# Interactive Identification of Granular Materials using Force Measurements

Abstract
The ability to identify granular materials facilitates the emergence of various new applications in robotics, ranging from cooking at home to truck loading at mining sites. However, granular material identification remains a challenging and underexplored area. In this work, we present a novel interactive material identification framework that enables robots to identify a wide range of granular materials using only a force-torque sensor for perception. Our framework, comprising interactive exploration, feature extraction, and classification stages, prioritizes simplicity and transparency for seamless integration into various manipulation pipelines. We evaluate the proposed approach through extensive experiments with a real-world dataset comprising 11 granular materials, which we also make publicly available. Additionally, we conducted a comprehensive qualitative analysis of the dataset to offer deeper insights into its nature, aiding future development. Our results show that the proposed method is capable of accurately identifying a wide range of granular materials solely relying on force measurements obtained from direct interaction with the materials. Code and dataset are available at: https://irobotics.aalto.fi/indentify_granular/.
I Introduction
The ability to accurately identify granular materials holds great potential across various domains of robotics, paving the way for future advancements in assistive robotics as well as autonomous working machines. In assistive robotics, being able to distinguish different food ingredients from each other is vital for planning and performing food-related downstream tasks such as assistive feeding or assistive cooking. Similarly, in industries like mining or construction, autonomous working machines must possess the capacity to differentiate between different granular materials for tasks such as truck loading and unloading or soil cultivating. Furthermore, material identification serves beyond mere manipulation tasks; it can also be employed for anomaly detection and quality assurance in production lines, contributing to overall efficiency and reliability.
Prior works on granular material identification rely either on tactile array sensors or audio sensing and identify materials based on their vibration ”signature” when the material is shaken in a closed container [eppe2018deep, jin2019open, chen2016learning, wang2023stev]. However, the lack of direct interaction means that the models’ classification decisions are based on material properties that may not best correspond to the material dynamics during manipulation, making the approaches suboptimal for use as a part of a manipulation pipeline.
To address this gap, we introduce a novel interactive perception framework that identifies granular materials by interacting with them directly. Specifically, our method measures feedback from the interaction with a standard 6-axis force-torque sensor (F/T sensor) and subsequently classifies the materials based on the features extracted from the raw F/T measurements. The F/T sensor is mounted on the robot’s end effector, which is partially submerged in the material during a predefined motion. We opt for a pure F/T sensor approach due to its simplicity and widespread availability of cost-effective sensors, ensuring ease of implementation across diverse real-world settings. The proposed approach enables robots to accurately identify the materials they interact with, enhancing their ability to understand and respond effectively to the environment. To the best of the authors’ knowledge, this work is the first attempt to identify granular materials using only a regular F/T sensor for perception. The contributions of this work include:
  • A novel interactive perception framework powered with a classification model tailored for granular material identification.
  • A novel real-world dataset of force-torque measurements containing 62 samples out of 11 different materials, totaling 682 individual samples.
  • A comprehensive analysis of the dataset, facilitates a deeper understanding of its underlying characteristics, the distinguishability of materials, and identification of their distinctive features.
II Related work
To put our work in context, we next review two complementary viewpoints, interactive perception, and granular material identification. Considering interactive perception, we position our focus on works that rely on 6-axis F/T sensors since they are closest related to our work.
II-A Interactive perception with 6-axis F/T sensor
The majority of works on interactive perception with F/T sensors focus on identifying surface materials [liu12] or estimating surface properties [le2021probabilistic] of rigid bodies. Typically, these studies involve attaching the F/T sensor to an artificial fingertip or similar end-effector, which interacts with the material surface to explore its properties. For example, Lam et al. [lam2014study] use this approach to identify 18 different surface materials. The authors reduce the raw data from the interaction into feature vectors of 3, 4, or 5 time-domain features and compare different classifiers, such as the k-nearest neighbor method, support vector machines, and neural networks, on their ability to distinguish the materials. Similarly, the work from Markert et al. [markert2021fingertip] proposed an interaction procedure for an artificial fingertip, which first gathers F/T data from 21 different surface materials, then classifies the materials with high accuracy using features derived from the F/T measurements’ spectral distributions. Furthermore, Le et al. [le2021probabilistic] integrated F/T sensing with visual information to estimate friction maps for multi-material rigid objects.
In addition, F/T sensors have also been used for object recognition and property estimation for rigid objects. For example, Higy et al. [higy2016combining] used 6-DoF F/T sensors combined with robot joint positions to perform haptic classification of rigid objects. However, information from multiple sensor modalities must be combined to achieve reasonably high classification accuracy, as poor performance is observed when each sensor operates independently. In [pavlic2023robotscale], the authors proposed a method that employs a 6-axis F/T sensor attached to the robot’s two-fingered end effector to estimate the mass and inertia of rigid objects.
Prior research highlights that F/T sensors can provide valuable information for various robotic perception tasks. Our work expands the scope of F/T sensing in robotics by introducing a novel application area focusing on deformable objects, in specific, granular materials. Additionally, we present an interpretable method for processing F/T sensor data from interaction with granular materials, further enhancing the applicability of F/T sensing in this domain.
II-B Granular Material Identification
The existing literature on granular material identification is sparse, with only a few initiatives focusing on identifying granular materials in closed containers. Some of those initiatives utilized audio recordings [eppe2018deep, jin2019open], while others leveraged vibration measurements [chen2016learning, wang2023stev]. In addition to identification, some works, such as [matl2020inferring] and [guo2023estimating], discuss the parameter estimation of granular materials.
In audio-based identification methods, a robot positions a closed container close to one or more microphones and then shakes the container. This shaking motion induces collisions among the granules inside the container and between the granules and the container walls, producing a sound pattern that varies according to material properties such as grain size and softness. For instance, Eppe et al. [eppe2018deep] used a humanoid robot to shake closed capsules filled with different materials, collected the audio recordings, and utilized a neural network to classify the samples. Although the study did not focus exclusively on granular materials, the majority of the ten materials examined can be considered granular. Similarly, Jin et al. [jin2019open] used a robot arm to rotate capsules filled with granular materials by 90 or 180 degrees and recorded audio samples from the interactions. The audio samples were then pre-processed by a noise reduction algorithm and transformed into the Mel-Frequency Cepstral Coefficient space to perform classification.
A major drawback with the audio-based methods is their proneness to get disturbed by background noise, as demonstrated in [eppe2018deep]. This might limit the approaches’ applicability in loud environments such as industrial production or construction sites. Furthermore, limited identification performance is reported for materials that do not make loud sounds when colliding with the container.
Vibration-based methods either complement data from microphones with other sensor modalities or use a completely different approach for vibration sensing. The work by Chen et al. [chen2016learning] is similar to audio-based methods since the work utilizes recordings from two microphones and a 3-axis accelerometer. The authors recorded data from 12 different materials that were sealed in containers that were shaken by a robot. However, the 12 materials differed significantly from each other, and only a few of them can be considered granular, with the rest being solid objects like an L-bracket or a piece of acrylic. Wang et al. [wang2023stev] showcased their novel electronic skin vibration sensor by classifying granular materials in a closed bottle. The authors used their soft robot gripper to shake the bottle and achieved high classification accuracy using different machine-learning models for the recorded signals, which were transformed from the time domain to the Fourier domain.
From a general viewpoint, we argue that the shaking-a-closed-container interaction is suboptimal when the classification information is used for manipulation-related downstream tasks. This is because the indirect interaction barely gives information about the material’s friction and flowing resistance properties, which are highly important for planning downstream manipulation actions such as scooping. Even though we are neither trying to estimate the friction parameters or flowing resistance explicitly, our proposed method directly interacts with the materials so that these properties are captured in the data. Thus, our method bases its classification decisions directly on the properties that affect the material’s manipulation-time dynamics the most.
III Interactive Material Identification
(a) F/T sensor and end-effector
(b) Starting position
Figure 1: Hardware setup and starting position
The proposed interactive material identification framework consists of three parts: interactive exploration, feature extraction, and classification. At the interactive exploration stage, the robot interacts with the materials directly and collects feedback with an F/T sensor. The feature extraction stage is used to process the raw F/T samples to enhance their classifiability. Finally, we employ a classification model based on error-correcting output codes (ECOC) and support vector machines (SVM) to identify the materials.
III-A Interactive Exploration
The interactive exploration stage involves the robot interacting with materials directly in an open container. The robot interacts with the materials with a cylinder-shaped end effector. The six-axis F/T sensor is attached between the robot’s wrist and the end effector, as depicted in Figure 1(a). In this setup, we utilize a 7-DoF Franka Emika Panda robot arm along with an ATI Mini45 F/T sensor. The F/T sensor measures data at a sampling frequency of 500 Hz throughout the motion.
To maintain a standardized exploration procedure, the height of the material surface and the robot starting position are predetermined and maintained consistently. The robot initiates its interaction with the material from the designated starting position, ensuring that the end effector is nearly fully submerged in the material, as depicted in Figure 1(b). We employ a Cartesian velocity controller to guarantee constant velocity regardless of the materials. The exploration trajectory entails linear motion along the x-axis at a velocity of 0.03 m/s for 3.5 seconds. The recording of F/T data is started a fraction of a second before the robot executes the exploration trajectory to ensure that the information-rich beginning of the interaction is fully recorded.
Although the F/T sensor in the robot records data for six axes, we focus only on the forces along the x-axis for classification purposes (the open dataset contains full data). This is because the movement of the robot’s end effector is limited to the x-axis direction, making the information measured along this axis the most significant. This decision simplifies the classification task by reducing the input dimensionality for the classification model.
III-B Feature extraction
To enhance the classifiability of the materials, we designed a feature space that takes into account their distinguishing characteristics, which are further discussed in Section IV. Specifically, we introduce a feature space representation called high-frequency magnitude histogram (HFMH) of the signal. In short, HFMH is the magnitude histogram of the F/T signal after high-pass filtering. We use a histogram with 100 bins evenly spaced in the range [-1.5, 1.5]. The range was defined to correspond to the maximum and minimum magnitudes present in the high-pass filtered F/T signals. For high-pass filtering, we use the Butterworth filter of order eight [proakisdigital], designed with respect to the desired cutoff frequency. Formally, we design a linear system defined by the transfer function
$H(z)=\frac{b_{0}+b_{1}z^{-1}+b_{2}z^{-2}+...+b_{8}z^{-8}}{1+a_{1}z^{-1}+a_{2}z^{-2}+...a_{8}z^{-8}}$ (1)
so that once the input signal is processed with the system, frequencies that are lower than the desired cutoff frequency get attenuated while higher frequencies are preserved. The best cutoff frequency, 23 Hz, was experimentally defined.
The designed feature space, neglecting the low-frequency content in the measurements, suggests that the materials are best separable based on the measurements’ high-frequency content, or lack thereof. The feature space also preserves information on the signal magnitudes to some extent, even though part of it is lost when the low frequencies are filtered out. The use of histograms leads to a compact representation with the cost of some information loss on the signal dynamics, that is, how the signal develops from the beginning to the end.
III-C Classification
We use the trinary ECOC model [escalera2008decoding] for our multiclass classification task. The ECOC model divides the multiclass classification problem into a set of binary classification tasks and includes a coding and decoding scheme to determine the final class for the input. For a $K$-class classification task, we have $L=K\cdot(K-1)/2$ binary learners. Formally, the ECOC decision rule is then defined as:
[…]
where $\hat{y}$ is the predicted class, $m_{kl}$ is an element of the trinary coding matrix $M\in\{-1,0,+1\}^{K\times L}$, and
$g(m_{kl},s_{l})=\max\{0,1-m_{kl}s_{l}\}/2$ (3)
is the binary loss. Variable $s_{l}$ stands for the binary classifier score for the input sample.
We use SVM with linear kernel functions as binary classifiers. Linear kernels were experimentally found to be most suitable for our task. They also allow a fair comparison when we compare the proposed feature space representation to baseline feature spaces.
The decision to use ECOC and SVMs with manually designed feature extraction instead of a data-driven approach, such as a classification neural network, is driven by two main reasons. Firstly, the available data is sparse, making it difficult to train a data-hungry neural network. Additionally, the time-intensive nature of real-world data collection makes it impractical to simply increase the dataset size. Secondly, one of the main purposes of material identification is to support downstream manipulation tasks. A transparent identification method combined with a hand-crafted feature extraction increases the feasibility of integrating knowledge from material identification into planning and control algorithms compared to black-box data-driven approaches. Furthermore, our approach enhances the understanding of the problem, materials, and measured F/T signals themselves, laying the groundwork for further advancements in the field.
IV Dataset and data qualitative analysis
IV-A Dataset
We employed the exploration procedure described in section III to collect a dataset of 62 individual samples from each of the 11 materials. These materials span a range of everyday items, including food ingredients, gardening materials, and animal care supplies. The complete list of the materials is shown in Figure 2. The materials were chosen to reflect diversity in density, grain size, grain shape, and surface properties of the grains. To prevent material compression during the data collection, the materials were stirred in the container after each measurement iteration. Furthermore, for materials that are susceptible to grain breakage from repeated interaction with the end effector (such as potting gravel, clay granules, cat litter, and oat flakes), fresh material was introduced during the data collection.
We introduce two variations of the dataset in our experiments: the easy dataset and the hard dataset. The easy dataset contains only 8 materials and is designed so that the majority of the materials are supposed to be distinct from each other. However, even the easy dataset contains seemingly similar materials, such as wheat flour and finely ground breadcrumbs. The hard dataset consists of 11 materials, including more materials with similar characteristics. For example, the hard dataset includes two different kinds of clay granules (gardening clay granules and cat litter), along with fine sugar as a new addition to the fine-grained materials. The hard dataset is designed to enhance the challenge of the identification task, further evaluating the effectiveness of the proposed method.
IV-B Qualitative data analysis
To gain deeper insights into the nature of the granular materials and how they differ, we conducted a comprehensive qualitative data analysis. The analysis revealed several differences, supporting our hypothesis that the proposed interactive identification framework can effectively differentiate between these materials. Figure 2 illustrates typical F/T measurements obtained from interactions with each material. It is evident that there are observable differences not only in the magnitude of the measurements but also in their high-frequency content. Furthermore, we can observe differences in the signal dynamics, that is, in the way the signals evolve from the initial samples toward the end of the signal. Although it is not possible to comprehensively analyze all materials separately within a limited space, in the following, we will discuss the distinctive features at a general level while highlighting the most interesting examples.
(a) Dry peas
(b) Rice
(c) Wheat flour
(d) Oat flakes
(e) Potting gravel
(f) Sunflower seeds
(g) Breadcrumbs
(h) Macaroni
(i) Fine sugar
(j) Cat litter
(k) Clay granules
Figure 2: Materials and examples of their typical F/T measurements (zoom in for a better view)
IV-B1 Signal magnitude
The most straightforward distinguishable factor between the different materials is the magnitude of the measurements, referred to as the signal magnitude. The signal magnitude indicates how much force is needed at each point in time during the interaction to maintain the constant end-effector velocity. Thus, the signal magnitude reflects information about the material’s flowing resistance, which correlates with its density and friction properties. Figure 2 highlights the differences between the materials. For instance, the average magnitude of high-density materials, like fine sugar, is significantly higher than that of lower-density materials, like wheat flour. However, relying solely on signal magnitude appears inadequate for identification, especially as the number of materials increases. This is evidenced, for example, by the similarities in the average signal magnitude of materials such as oat flakes and breadcrumbs.
IV-B2 Frequency content
Figure 2 also highlights distinct frequency-related patterns exhibited by various materials. For example, rice yields low-amplitude, rapid zig-zag patterns, whereas wheat flour demonstrates smoother behavior. Fine sugar shows consistent and relatively smooth, high-amplitude wave patterns. Macaroni, on the other hand, produces measurements with irregular and sharp turns. As observed, materials with larger grain sizes or high friction, stemming from irregular grain shapes or surface properties, tend to produce sharper and more irregular patterns in the measurements. We hypothesize these patterns are associated with the grain size and shape, as well as the slide and rolling friction of the grains. These properties determine how smoothly the robot end-effector can traverse the material.
Although the patterns affect the signals’ frequency domain representations, pure frequency domain signals are noisy and lack distinguishable patterns. We will also show in the section V that classification results using frequency domain representations are poor.
IV-B3 Signal dynamics
Signal dynamics refer to the qualitative patterns of how the signals develop from the beginning to the end. The signals, in general, exhibit non-stationarity, with the majority showing an increasing trend. This is expected, since the end-effector pushes the material forward during movement, causing the material to compress between the end-effector and the back edge of the container. Thus, a higher force is needed towards the end of the motion trajectory to maintain a constant velocity.
Some of the materials also display unique dynamics. For example, rice produces an easily identifiable high peak at the beginning of the signal, after which the signal descends close to zero and starts to gradually increase again towards the end. Another interesting, unique pattern can also be seen for clay granules. For clay granules, after an initial force increase, the readings take negative values. We associate this behavior with the dynamics of the material flow as the material moves away from the end-effector’s path and returns to fill the groove.
In summary, our analysis demonstrates that direct interaction with various granular materials yields meaningful force measurements. Distinct characteristics emerge for different materials in both the time and frequency domain. The provided analysis gives valuable insights for designing high-quality feature spaces that not only facilitate the classification of granular materials but also enable the estimation of more complex material properties, such as friction coefficients or flow dynamics, in the future. Based on our analysis, we hypothesize that a high-quality feature space for accurately distinguishing granular materials must consider the aforementioned characteristics of the measurements, including signal magnitude, frequency content, and signal dynamics.
V Experiments and Results